# A Proxy-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

### v1

@@ Appendix A Supplementary Materials
This supplementary material is organized as follows. We first discuss the attack success rate and benign accuracy under poisoned rates with and on the CIFAR-10 dataset in Sec. A.1 and Sec. A.2, respectively. Additionally, we presents the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec. A.3. Besides, Sec. A.4 presents the experimental results on the Tiny-ImageNet dataset. At last, Sec. A.5 shows the experiments on different pre-trained feature extractor .
A.1 Attack Success Rate Under Additional Poisoned Rates on the CIFAR-10 Dataset
We evaluate the attack success rate (ASR) under different poisoned rates on the CIFAR-10 dataset. Tables 5 and 6 present the ASR with trigger BadNets and Blended at poisoned rates of and , respectively. Results under poisoned rates of and are similar to those at .
(i) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.030 and 0.055 on BadNets and Blended triggers, respectively. The proxy-free method also achieves better attack results than FUS in most cases (27/36). The combination of these two methods achieves superior results in all 36 settings.
(ii) Experiments under poisoned rates of : Our proxy-free approach outperforms random selection in all cases, with average boosts of 0.023 and 0.04 on BadNets and Blended triggers, respectively. The proxy-free method also achieves better attack results than FUS in most cases (26/36). The combination of these two methods achieves superior results in all 36 settings.
@@ A.1 Attack Success Rate Under Additional Poisoned Rates on the CIFAR-10 Dataset
A.2 Benign Accuracy Under Additional Poisoned Rates on the CIFAR-10 Dataset
We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoned rates on the CIFAR-10 dataset. Table 7 and Table 8 illustrate the benign accuracy with trigger BadNets and Blended under poisoned rates of and , respectively. The results show the similar BA with four strategies, which demonstrates that our proposed strategy has no negative impact on benign accuracy.
@@ A.2 Benign Accuracy Under Additional Poisoned Rates on the CIFAR-10 Dataset
A.3 Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
This section presents the attack success rate (ASR) under additional target classes on the CIFAR-10 dataset. Table 9 illustrates the attack success rate with trigger BadNets and Blended under target category 3. Compared with random selection, our proxy-free approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers. The proxy-free method also achieves better attack results in most cases compared to FUS (22/24). The combination of these two achieves superior results in all 24 settings. These outcomes consistently confirm the conclusions we reached in the paper.
@@ A.3 Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
A.4 Experiments on Tiny-ImageNet Dataset
In Table 10, it can be observed that our method outperforms the random method in terms of attack success rates of poisoned samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively. In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases. When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method. These findings provide strong evidence supporting the effectiveness of our proposed method.
@@ A.4 Experiments on Tiny-ImageNet Dataset
A.5 Experiments on Different Pre-trained Feature Extractor

### v2

@@ Appendix A Supplementary Materials
This supplementary material is organized as follows. We first discuss the attack success rate and benign accuracy under poisoning rates with and on the CIFAR-10 dataset in Sec. A-B and Sec. A-C, respectively. Additionally, we present the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec. A-D. Besides, Sec. V-C presents the experimental results on the Tiny-ImageNet dataset.
A-A The Proof of Theorem 1
Theorem 2.
Suppose the training dataset consists of benign samples and poisoned samples , whose images are i.i.d. sampled from uniform distribution and belonging to classes. Assume that the DNN is a multivariate kernel regression and is trained via . For the expected predictive confidences over the target label , we have: , where is poisoned testing samples of attacks.
Proof 1 (Proof of Theorem 1:).
We treat our model as a m-way kernel least square classifier and use a cross-entropy loss for training the kernel. The output of is a m-dimensional vector. Let us assume be expected predictive confidences corresponding to the target class . Following previous works [15, 16], we know the kernel regression solution for neural tangent kernel (NTK) [22] is:
where is the RBF kernel and . Since the training samples are evenly distributed, there are benign samples belonging to . Without loss of generality, we assume the target label while others are . Then the regression solution can be re-formulated to:
Therefore, following [30], we have:
Similar to [16], we also remove the term because typically don’t belong to the target and , otherwise the attacker has no incentive to craft poisoned samples.
When close to , which implies that the poisoning rate close to , the attacker can achieve the optimal attack efficacy [14, 27]. Given , we have:
we here let in , because typically be similar as and . Therefore, we have:
A-B Attack Success Rate Under Additional poisoning rates on the CIFAR-10 Dataset
We evaluate the attack success rate (ASR) under different poisoning rates on the CIFAR-10 dataset. Tables VI and VII present the ASR with trigger BadNets and Blended at poisoning rates of and , respectively. Results under poisoning rates of and are similar to those at .
(i) Experiments under poisoning rates of : Our PFS approach outperforms random selection in all cases, with average boosts of 0.030 and 0.055 on BadNets and Blended triggers, respectively. The PFS method also achieves better attack results than FUS in most cases (27/36). The combination of these two methods achieves superior results in all 36 settings.
(ii) Experiments under poisoning rates of : Our PFS approach outperforms random selection in all cases, with average boosts of 0.023 and 0.04 on BadNets and Blended triggers, respectively. The PFS method also achieves better attack results than FUS in most cases (26/36). The combination of these two methods achieves superior results in all 36 settings.
@@ A-B Attack Success Rate Under Additional poisoning rates on the CIFAR-10 Dataset
A-C Benign Accuracy Under Additional poisoning rates on the CIFAR-10 Dataset
We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoning rates on the CIFAR-10 dataset. Table VIII and Table IX illustrate the benign accuracy with trigger BadNets and Blended under poisoning rates of and , respectively. The results show the similar BA with four strategies, which demonstrates that our proposed strategy has no negative impact on benign accuracy.
@@ A-C Benign Accuracy Under Additional poisoning rates on the CIFAR-10 Dataset
A-D Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
This section presents the attack success rate (ASR) under additional target classes on the CIFAR-10 dataset. Table X illustrates the attack success rate with trigger BadNets and Blended under target category 3. Compared with random selection, our PFS approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers. The PFS method also achieves better attack results in most cases compared to FUS (22/24). The combination of these two achieves superior results in all 24 settings. These outcomes consistently confirm the conclusions we reached in the paper.
@@ A-D Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
A-E Experiments on CIFAR-100 Dataset
The results for CIFAR-100 in Table XI are similar to those for CIFAR-10. Our PFS approach outperforms random selection in all cases, with average boosts of 0.047, 0.068, and 0.037 for BadNets, Blended, and Optimized triggers, respectively. Compared to FUS, PFS achieves better attack results in 49 out of 54 settings.
@@ A-E Experiments on CIFAR-100 Dataset
A-F Experiments on Different Pre-trained Feature Extractor